OpenStudy (anonymous):

Evaluate the integral:

OpenStudy (anonymous):

\[\int\limits_{}^{}\csc^{6}2x dx\]

OpenStudy (anonymous):

HINT: \[\int\limits(\csc ^{n}u) du = \frac{ -1 }{ n-1 }\cot u (\csc ^{n-2}u) + \frac{ n-2 }{ n-1 }\int\limits(\csc ^{n-2}u)du\]